OpenStudy (anonymous):

Simplify these rational expressions & find the remainders x^4+5x^3-2x^2+1/x^2+1, & 3x^3+4x+5/x+3

OpenStudy (anonymous):

the remainder of x^4+5x^3-2x^2+1/x^2+1 =( a-5x-2, b=5x-4, c=-5x-2, d= -5x-4) the remainder of 3x^3+4x+5/x+3= (a= 0, b=5,c=-88, d=93)